12. Joe can row a boat 3 mph faster than he can swim. He is able to row 10 miles in the same time it takes him to swim 4 miles. Find Joes rate of swimming and rowing
Answer #1

let joe's rate of swimming = x mph

rate of rowing = x + 3

applying distance rate time formula

10 / ( x+ 3) = 4 / x

10x = 4 ( x+ 3)

10x = 4x + 12

6x = 12

x = 2

hence , rate of swimming = 2 mph

and rate of rowing = 5 mph